if 1.82g of fluorine-18 is shipped at 5:00pm, how many milligrams of the radioisotope are still active if the sample arrives at the nuclear medicine laboratory at 9:30am the next morning
Answer #1

Fluorine-18 has a half-life of 110 min therefore rate of reaction is given as

k = 0.693 / t1/2 = 0.693 / 110 min = 6.3 x 10^-3 /min

and time 5;00 pm to 9;30 am ; 16 hours    30 minutes

or 990 minutes

[F0] = 1.82 g

using the following equation we get

k = 2.303 / t * ln [F0] / [Ft]

6.3 x 10^-3 /min = 2.303/ 990 min ln [F0] / [Ft]

2.71 = ln [F0] -ln [Ft]

2.71= ln 1.82 - ln [Ft]

2.71- 0.599 = -ln [Ft]

[Ft]= 10^-2.11

= 0.121